Find the measure of an angle if its measure is 26° more than three times its complement. Help! Thanks (:
Answer by algebrahouse.com(1659)

x = the angle
90 - x = the complement {complementary angles add up to 90°}

x = 3(90 - x) + 26 {the angle is equal to 26 more than three times its complement}
x = 270 - 3x + 26 {used distributive property}
x = -3x + 296 {combined like terms}
4x = 296 {added 3x to each side}
x = 74 {divided each side by 4}

The angle is 74°

For more help from me, visit: www.algebrahouse.com